Project

General

Profile

Ops/Comm Scheduling meeting : 22/8/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
  • Status of alarm clearance (DC/Ajit)
  • Hydrogen system update and plans (JBoehm)
    • Update to search procedure documentation for hydrogen system (Craig/Henry)
  • Cycle 2017 Issues - revisit after the MICE-ISIS LH2 sign-off
  • Hall/Operations plan : 17-08-17-CW.xlsx
  • Operations startup
    • Initial discussion on system readiness : Wednesday August 23rd in CR25 in R1 @ 11
    • Pre-startup meeting : Monday Sep 4th
    • Pulser run : Sep 7th/8th
    • Beam tuning and beam test run : sometime in the week starting Sep 11th. Will ask ISIS for 8 hours during this week. ISIS scheduling meeting is on August 24th.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    Yes!
    No :(
  • AOB

Minutes

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
    Missed planned LH2 walk-about. The pipework components are now done and are being installed; will then cut the final piece to length and fit end today/first thing tomorrow. Shed on roof almost complete: exVenturis in at the moment. Phil is down working on the software items; will check with Omron before finishing the network lockout. If the pipework is completed, do the leak test tomorrow p.m. and then close up gas panel and affix ATEX signage.
    Last Friday checked the SS voltage-to-zero issue: looks to be in software as the relevant control sequence exists but is not getting triggered. Also discovered that the SSU M2 QPS channel has no signal coming into it. Need to get a drawing so that Craig and Phil can check the cabling.
    Test the Target Frame later in the week.
    A/C unit 4 now also has an alarm, about changing the filter.
    Paul K and Melissa are practising re-calibrating the Trackers. Cooling performance has improved after pump-purge, but more dead channels.
    Need to lock Hall up at some point to re-test the QPS system, but also to open DSA to check Proton Absorber.
  • Status of alarm clearance (DC/Ajit)
    Have found a number of issues which are just changes to values in the spreadsheet - will do these en masse.
    A number of environment variables are stuck: possibly a bug in EPICS or the IOC (invalidity of low-level PV isn't cascading upward). Will be cleared when IOC restarted but need something to monitor them.
  • Hydrogen system update and plans (JBoehm)
    Low pressure alarm on SS compressor - slow leak.
    Need the adsorbers within next week to have time to do work.
    N.B. need a Helium bottle at West end of LH2 Mezz, as won't be able to get them past Gas Panel once LH2 system live.
  • Update to search procedure documentation for hydrogen system (Craig/Henry) - in progress
  • Operations startup
    • Initial discussion on system readiness : Wednesday August 23rd in CR25 in R1 @ 11
    • Pre-startup meeting : Monday Sep 4th
    • Pulser run : Sep 7th/8th
    • Beam tuning and beam test run : sometime in the week starting Sep 11th. Will ask ISIS for 8 hours during this week. ISIS scheduling meeting is on August 24th.
      Someone should attend in person - CRogers got volunteered.
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB

17-08-17-CW.xlsx (50.5 KB) Boyd, Steven, 21 August 2017 12:27